OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (hunus):

Do you have any idea where to start?

OpenStudy (texaschic101):

x - 2y = 2 --> x = 2y + 2 now sub 2y + 2 in for x in the other equation 3x + y = 6 3(2y + 2) = 6 6y + 6 = 6 6y = 6 - 6 6y = 0 y = 0/6 = 0 now sub 0 in for y x - 2y = 2 x - 2(0) = 2 x - 0 = 2 x = 2 check... 3x + y = 6 3(2) + 0 = 6 6 + 0 = 6 6 = 6 (correct) ANSWER : (2,0)

OpenStudy (texaschic101):

Do you have any questions...if so, please ask them

OpenStudy (texaschic101):

This is the substitution method....it can also be solved by using the elimination method

OpenStudy (anonymous):

3x + y = 6 3(2y + 2) = 6 <--- should be 3(2y+2) + y = 6 # the '+y' is gone. Otherwise, it is good.

OpenStudy (texaschic101):

I messes up.....let me do it again x - 2y = 2 --> x = 2y + 2 now sub 2y + 2 in for x in the other equation 3x + y = 6 3(2y + 2) + y = 6 (distribute through the parenthesis) 6y + 6 + y = 6 (combine like terms) 7y + 6 = 6 7y = 6 - 6 7y = 0 y = 0/7 = 0 x is solved the same way ...in the above answer I did. It doesn't make a difference in the answer because y is still 0. thanks @franciscanmonk for finding my mistake :)

OpenStudy (anonymous):

No probs. I didn't really check if it would make a difference though. haha.

OpenStudy (texaschic101):

I kinda freaked out....I thought I just gave the wrong answer....when I re-did it, I found that it wouldn't have made a difference.

OpenStudy (anonymous):

It's all good. Don't worry :)

OpenStudy (texaschic101):

That is, it wouldn't have made a difference in the answer. But if you had to show your work, that would be a problem.....

OpenStudy (anonymous):

You're right. :)